9900 sec = how many hours?
Found 2 solutions by math_tutor2020, mananth:
Answer by math_tutor2020(3816)


Answer: 2.75 hours (exact)

Work Shown

The units "sec" and "min" cancel out based on how the conversion fractions are set up.

Or a slight shortcut

The shortcut works because 1 min = 60 sec and 1 hr = 60 min, so 1 hr = 60*60 = 3600 sec

Answer by mananth(16946)

9900 sec = how many hours?
1 hour = 60 minutes
1minute = 60 secs
there fore 60*60 = 3600 seconds = 1 hour
9900 seconds = 9900/3600=2.75 hours